OpenStudy (anonymous):

Can someone check this answer for me? Which of the following is equivalent to i14? A.)i B.)-i C.)-1 D.)1

OpenStudy (anonymous):

i got B

OpenStudy (anonymous):

what was your answer?

OpenStudy (anonymous):

what does i14 mean?

OpenStudy (anonymous):

i think its the equivalent

OpenStudy (anonymous):

If it is simply i14, the answer is still i since you're just simply multiplying. If you mean \[i ^{14}\] the answer is B.

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

You're welcome. Since you already have this choices, once you encounter this problem again always do this. Simply divide the exponent of the i by 4 and if the remainder is 1 then answer is i; remainder is 2 answer is -1 remainder is 3, answer is -i remainder is 0, answer is 1.

OpenStudy (anonymous):

the answer must be C = -1 because i^14=i^2= -1

OpenStudy (anonymous):

wait is it c or b?

OpenStudy (anonymous):

it ur question had i^15 then it would have been equal to i^3 which equals to -i. But in this case answer is B , IF i stands for iota.

OpenStudy (anonymous):

pls fan me if u like my help...thanx

OpenStudy (anonymous):

gud luck with ur maths

OpenStudy (anonymous):

i stands for imaginary number. i.e sqrt(-1) I dont know what iota is.

OpenStudy (anonymous):

iota means a very small quantity as far as i understand

OpenStudy (anonymous):

iota = sqrt(-1).

OpenStudy (anonymous):

the one which u r talking about that is EPSILON

OpenStudy (anonymous):

IOTA cant be a small quantity becoz its imaginary

OpenStudy (anonymous):

katie, \[i ^{1}= i, i ^{2}=-1 , i ^{3}= -i, i ^{4}=1\]

OpenStudy (anonymous):

from wikipedia, In mathematics and statistics, a unit vector is often denoted by a lowercase iota letter with a superscript caret: .

OpenStudy (anonymous):

Sorry. it is C

OpenStudy (anonymous):

Hey...congrats...now we have a concensus on the answer...good one...

OpenStudy (anonymous):

from thefreedictionary, iota - a tiny or scarcely detectable amount

OpenStudy (anonymous):

yes, the answer is C.